# Hawking–Page transition on a spin chain

### v1

@@ Abstract
[…]
@@
In this letter, we examine another important aspect of gravity: the formation of a black hole (BH). Unlike wormholes, which harbour innate quantum properties, BHs are prominently thermal entities. Hence, they can be analysed through conventional means such as finite-temperature lattices. A BH forms in anti-de Sitter (AdS) spacetime with the Hawking–Page transition Hawking and Page (1983). Our main point asserts that this phenomenon can be simulated on a simple spin chain, the Heisenberg XX chain.
[…]
@@
[…]
@@
where $S$ is the entropy and $s_{\text{low}},s_{\text{high}}$ are $\mathcal{O}(1)$ constants. Enormous effort has been put to test and sharpen this prediction Sundborg (2000); Polyakov (2002); Balasubramanian et al. (2002); Aharony et al. (2004); Liu (2004); Berenstein (2004); Alvarez-Gaume et al. (2005); Festuccia and Liu (2006); Basu and Wadia (2006); Alvarez-Gaume et al. (2006); Biswas et al. (2007); Chang and Yin (2013); Hosseini et al. (2017); Cabo-Bizet et al. (2019a); Choi et al. (2018a, b); Benini and Milan (2020a, b); Honda (2019); Arabi Ardehali (2019); Cabo-Bizet et al. (2019b); González Lezcano and Pando Zayas (2020); Larsen et al. (2020); Cabo-Bizet and Murthy (2020); Arabi Ardehali et al. (2020); Benini et al. (2020); David et al. (2020); Cabo-Bizet et al. (2020); Agarwal et al. (2021); González Lezcano et al. (2021); Copetti et al. (2022); Goldstein et al. (2021); Amariti et al. (2021a, b); Cassani and Komargodski (2021); Aharony et al. (2021); Uhlemann (2021); Hong (2021); Ezroura et al. (2022); Imamura (2021); Cabo-Bizet (2022); Murthy (2023); Boruch et al. (2022); Honda and Yoda (2022); Lee (2022); Huang (2022); Imamura (2022); Holguin and Wang (2022); Eleftheriou (2023); Choi et al. (2023); Chang and Lin (2023); Chen et al. (2023a); Liu and Rajappa (2023); Lin (2023); Eniceicu (2023); Beccaria and Cabo-Bizet (2023a); Chang et al. (2023a); Chen et al. (2023b); Ekhammar et al. (2023); Bigazzi et al. (2023); Beccaria and Cabo-Bizet (2023b); Arabi Ardehali et al. (2023); Chang et al. (2023b); Eleftheriou et al. (2023).
@@ I Hawking–Page transition
[…]
@@
$e^{\hat{S}}=\int_{0}^{\infty}\sigma\mathrm{d}\sigma\oint\mathrm{d}U\leavevmode\nobreak\ \exp\left[-N^{2}\left(\frac{\sigma^{2}}{{4a}}-\mathcal{F}(\sigma)\right)\right],$ (9)
@@ II Mapping to the spin chain
[…]
@@
Note that $\tilde{T}$ is the temperature of the chain, unrelated to the BH temperature $T$. Furthermore, (16) only depends on $J\equiv\tilde{J}/\tilde{T}$. Eq. (15) is not normalized and should be divided by $\mathrm{Tr}\left(e^{-H_{\text{XX}}/\tilde{T}}\right)$. It is more convenient to work with the ratios
@@
With this prescription, ^$\hat{\mathcal{L}}_{N}$ remains finite when $L\to\infty$.
@@
We thus propose the following setup. ^$\hat{\mathcal{L}}_{N}(J)$ is measured in the regime $1\ll N\ll L$, with $J$ as above. The experiment is run several times, sampling $J$ from a random Gaussian distribution with standard deviation $2a$. The average echo is computed:
^$\left\langle\sqrt{\hat{\mathcal{L}}_{N}}\right\rangle_{2a}=\int_{0}^{\infty}J\mathrm{d}J\leavevmode\nobreak\ e^{-\frac{J^{2}}{4a}}\sqrt{\hat{\mathcal{L}}_{N}(J)}$ (25)
The dependence on $\tilde{T}$ is reabsorbed in the integration variable, thus the only free parameter is $a>0$. The measurement can be performed at arbitrary non-zero temperature $\tilde{T}$.
@@
The symbol $\sim$ means equality up details negligible when $N\gg 1$. The temperature $\tilde{T}$ is unimportant and distinct from the BH temperature $T$, which enters (26) via $a$. Moreover, (26) is expressed using ^$\sqrt{\hat{\mathcal{L}}_{N}}$, rather than $\hat{\mathcal{G}}_{N}$, because only the former is a measurable quantity.
@@ III Refined probes
We explain four ways to refine (26) for more accurate tests of the Hawking–Page transition. Details are in §B.
@@
2. The spin chain can reproduce the perturbative corrections in the ’t Hooft coupling to the BH entropy (§B.3). Surprisingly, this major conceptual step requires no further sophistication of the experimental setup. These corrections are incorporated without complicating the Hamiltonian nor the initial state. The implementation simply requires that the Gaussian random coupling is sampled with random standard deviation — see §B.3 for details. The BH entropy is computed by ‘nested’ averages of the Loschmidt echo.
3. A fine probe of the transition, from the CFT side of the holographic duality, is the Polyakov loop. On general grounds, its expectation value $\mathcal{P}$ vanishes if $T<T_{\mathrm{HP}}$ and is non-vanishing in a BH phase $T>T_{\mathrm{HP}}$. Therefore, the study of $\mathcal{P}$ tests the properties of the transition beyond the jump in entropy. In the spin chain, the Polyakov loop is easily realized as an impurity in the state $\lvert\psi_{0}\rangle$ Pérez-García et al. (2022). We elaborate on this aspect in §B.4.
@@ IV Discussion
[…]
Unlike the planar integrability of $\mathcal{N}=4$ SYM (§C), the XX chain used here is an auxiliary device manifesting the phase transition, rather than being a subsector of SYM. We map one specific observable to the averaged thermal Loschmidt echo in the spin chain.
The outcome highlights a universality property: introducing entropy in the 1d chain through the average, it emulates the behaviour of (superficially unrelated) BH systems in (4+1)d.
It is natural to inquire about the extent to which the spin chain captures structural properties of BH physics, besides the jump in entropy at the Hawking–Page transition. In §D, we elaborate on two concrete avenues for future research, aimed at characterizing the BH microstates.
@@ Acknowledgements.
We are grateful Hamed Adami, Elliott Gesteau and Sofyan Iblisdir for discussions. LS thanks the Departamento de Análisis Matemático y Matemática Aplicada, Universidad Complutense de Madrid for warm hospitality at various stages of this project. The work has been financially supported by Comunidad de Madrid (grant QUITEMAD-CM, ref. P2018/TCS-4342), by Universidad Complutense de Madrid (grant FEI-EU-22-06), by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2025 Agenda. DPG also acknowledges financial support from the Spanish Ministry of Science and Innovation (“Severo Ochoa Programme for Centres of Excellence in R&D” CEX2019- 000904-S and grant PID2020-113523GB-I00) as well as from CSIC Quantum Technologies Platform PTI-001. The work of LS is supported by the Shuimu Scholars program of Tsinghua University and by the Beijing Natural Science Foundation project IS23008 “Exact results in algebraic geometry from supersymmetric field theory”.
@@
[…]
@@ A.2 Thermal partition function
[…]
@@
$S=\log(\text{\# miscrostates}).$ (31)
[…]
@@ A.4 Weak coupling corrections and deconfinement
In summary, the simplified partition function (8) captures all the distinctive properties of the Hawking–Page transition, and a more realistic approximation, that includes a coexistence region for thermal AdS and BH, is described by the improved formula (32).
@@
B.1 Improving the estimate of the BH entropy

### v2

@@ Abstract
The accessibility of the Hawking–Page transition in AdS5 through a 1d Heisenberg spin chain is demonstrated. We use the random matrix formulation of the Loschmidt echo for a set of spin chains, and randomize the ferromagnetic spin interaction. It is shown that the thermal Loschmidt echo, when averaged, detects the predicted increase in entropy across the Hawking–Page transition. This suggests that a 1d spin chain exhibits characteristics of black hole physics in 4+1 dimensions. We show that this approach is equally applicable to free fermion systems with a general dispersion relation.
@@
In this paper, we examine another important aspect of gravity: the formation of a black hole (BH). Unlike wormholes, which harbour innate quantum properties, BHs are prominently thermal entities. Hence, they can be analysed through conventional means such as finite-temperature lattices. A BH forms in anti-de Sitter (AdS) spacetime with the Hawking–Page transition Hawking and Page (1983). Our main point asserts that this phenomenon can be simulated on a simple spin chain, the Heisenberg XX chain.
To elaborate, we shall begin with thermal AdS5. As the temperature $T$ rises, a BH solution emerges, which becomes dominant above a threshold $T_{\mathrm{HP}}$ Hawking and Page (1983). It can be inferred that, at $T>T_{\mathrm{HP}}$, the entropy would be significantly larger than thermal AdS, due to the presence of the BH horizon. This follows from the Bekenstein–Hawking formula Bekenstein (1973); Hawking (1975)
@@
with $A_{\mathrm{hor}}$ the area of the BH horizon and $G_{\text{N}}$ the Newton constant in AdS5 (in units $c=\hbar=1$).
The celebrated AdS/CFT correspondence Maldacena (1998); Gubser et al. (1998); Witten (1998a) posits that gravitational asymptotically-AdS solutions are encoded into a conformal field theory (CFT) on the boundary. The most widely studied example, and the one we will focus on, is the duality between AdS5 and 4d $\mathcal{N}=4$ super-Yang–Mills theory (SYM) Maldacena (1998).
@@
where $S$ is the entropy and $s_{\text{low}},s_{\text{high}}$ are $\mathcal{O}(1)$ constants. Enormous effort has been put to test and sharpen this prediction Sundborg (2000); Polyakov (2002); Balasubramanian et al. (2002); Aharony et al. (2004); Liu (2004); Berenstein (2004); Alvarez-Gaume et al. (2005); Festuccia and Liu (2006); Basu and Wadia (2006); Alvarez-Gaume et al. (2006); Biswas et al. (2007); Chang and Yin (2013); Hosseini et al. (2017); Cabo-Bizet et al. (2019a); Choi et al. (2018a, b); Benini and Milan (2020a, b); Honda (2019); Arabi Ardehali (2019); Cabo-Bizet et al. (2019b); González Lezcano and Pando Zayas (2020); Larsen et al. (2020); Cabo-Bizet and Murthy (2020); Arabi Ardehali et al. (2020); Benini et al. (2020); David et al. (2020); Cabo-Bizet et al. (2020); Agarwal et al. (2021); González Lezcano et al. (2021); Copetti et al. (2022); Goldstein et al. (2021); Amariti et al. (2021a, b); Cassani and Komargodski (2021); Aharony et al. (2021); Uhlemann (2021); Hong (2021); Ezroura et al. (2022); Imamura (2021); Gaiotto and Lee (2021); Cabo-Bizet (2022); Murthy (2023); Boruch et al. (2022); Honda and Yoda (2022); Lee (2022); Huang (2022); Imamura (2022); Holguin and Wang (2022); Eleftheriou (2023); Choi et al. (2023); Chang and Lin (2023); Chen et al. (2023a); Liu and Rajappa (2023); Lin (2023); Eniceicu (2023); Beccaria and Cabo-Bizet (2023); Chang et al. (2023a); Chen et al. (2023b); Ekhammar et al. (2023); Bigazzi et al. (2023); Beccaria and Cabo-Bizet (2024); Arabi Ardehali et al. (2023); Chang et al. (2023b); Lee (2024); Eleftheriou et al. (2023).
@@ I Hawking–Page transition
In this paper we focus on the BH formation in maximally-supersymmetric AdS5, whose holographic dual is $\mathcal{N}=4$ SYM. On the one hand, maximal supersymmetry plus conformal symmetry tightly constrain the theory, making it the perfect playground to explore 4d physics. On the other hand, it bears universal properties akin to pure Yang–Mills and QCD Polchinski and Strassler (2002); Bern et al. (2004); Caron-Huot and Moore (2008), befitting phenomenological applications.
The AdS/CFT correspondence dictates that the Hawking–Page transition Hawking and Page (1983) in AdS5 is dual to a first order transition in $\mathcal{N}=4$ SYM on Euclidean $\mathbb{S}^{3}\times\mathbb{S}^{1}$ Witten (1998b) (details in §A), with thermal $\mathbb{S}^{1}$ of radius $1/T$. An effective description for this theory, in the weak coupling limit, was devised in Sundborg (2000); Polyakov (2002); Aharony et al. (2004). The result, written in terms of the AdS5 entropy $S$, is:
@@
$e^{\hat{S}}=\int_{0}^{\infty}\sigma\mathrm{d}\sigma\leavevmode\nobreak\ \exp\left[-N^{2}\left(\frac{\sigma^{2}}{{4a}}-\mathcal{F}(\sigma)\right)\right],$ (9)
@@ II Mapping to the spin chain
Our novel claim is that the Hawking–Page transition in supersymmetric AdS5 can be encoded in a 1d quantum spin chain. The simulation of such thermodynamic process requires a chain at finite temperature. The quantity detecting the sudden change in entropy is the Loschmidt echo Gorin et al. (2006), an observable tailored to quantify how chaotic a system is. It has already found broad application in BH physics (e.g. del Campo et al. (2018); Chenu et al. (2018, 2019); Almheiri and Lin (2022)). Our insight is to interpret (9) as a coupling average of the Loschmidt echo in the XX chain.
@@
Note that $\tilde{T}$ is the temperature of the chain, unrelated to the BH temperature $T$. Furthermore, (16) only depends on $J\equiv\tilde{J}/\tilde{T}$. In (15) we used the standard replacement $it\mapsto 1/\tilde{T}$. In this way, $\mathcal{L}_{N}$ is not a probability, because it is not canonically normalized. It is thus convenient to work with the ratios
@@
One advantage of this prescription, is that ^$\hat{\mathcal{L}}_{N}$ remains finite when $L\to\infty$.
@@
We thus propose the following setup. ^$\hat{\mathcal{L}}_{N}(J)$ is measured in the regime $1\ll N\ll L$, with $J$ as above. The experiment is run several times, sampling $J$ from a suitably chosen random distribution. The average echo is computed:
^$\left\langle\sqrt{\hat{\mathcal{L}}_{N}}\right\rangle_{2a}=\int_{0}^{\infty}\frac{J\mathrm{d}J}{2a}\leavevmode\nobreak\ e^{-\frac{J^{2}}{4a}}\sqrt{\hat{\mathcal{L}}_{N}(J)}$ (25)
The dependence on $\tilde{T}$ is reabsorbed in the integration variable, thus the only free parameter is $a>0$. We refer to $2a$ as the standard deviation. The measurement can be performed at arbitrary non-zero temperature $\tilde{T}$.
@@
The symbol $\sim$ means equality up to details negligible when $N\gg 1$. The temperature $\tilde{T}$ is unimportant and distinct from the BH temperature $T$, which enters the left-hand side of (26) via $a$. Moreover, (26) is expressed using ^$\sqrt{\hat{\mathcal{L}}_{N}}$, rather than $\hat{\mathcal{G}}_{N}$, because only the former is a measurable quantity.
@@
Interpretation of the result
The AdS5/CFT4 correspondence led us to an equality relating the large $N$ BH entropy and the coupling averaged Loschmidt echo. On the BH side of (26), the only parameter is the temperature $T$, which appears through $a$. The dependence (7) is monotonic, thus it can be inverted and the BH temperature can be thought of as a function of $a$.
On the spin chain side of (26), the free parameter $a$ determines how the coupling is sampled. It can be varied freely, and does not require tuning. The experiment is repeated many times, varying $J$ as sampled randomly according to (25), and computing the average of the measured Loschmidt echo. The right-hand side of (26) is a controlled setup that should experience a sharp change in behaviour, when the procedure is performed sampling the coupling at $a<1$ versus $a>1$ (Fig. 1).
In conclusion, our formula indicates how to concretely test and validate the Hawking–Page prediction, in the spirit of quantum simulation.
@@ III Refined probes
We explain four ways to refine (26) for more accurate tests of the Hawking–Page transition. We discuss:
  1. 1.
    The analysis of the order parameters for the transition;
  2. 2.
    The inclusion of corrections by higher operators;
  3. 3.
    The inclusion of corrections in the ’t Hooft coupling;
  4. 4.
    The case of complex fugacities for the black hole charges.
The technical details are relegated to §B.
Order parameters
1. A fine probe of the transition, from the CFT side of the holographic duality, is the Polyakov loop. This is defined as a Wilson loop winding around the Euclidean time direction. On general grounds, its expectation value $\mathcal{P}$ vanishes if $T<T_{\mathrm{HP}}$ and is non-vanishing in a BH phase $T>T_{\mathrm{HP}}$. Therefore, the study of $\mathcal{P}$ tests the properties of the transition beyond the jump in entropy.
We compute $\mathcal{P}$ in the matrix model and find that indeed it is an order parameter for the phase transition (see §B.1 for the derivation).
In the spin chain, the Polyakov loop is realized as an impurity in the preparation of the initial state. That is, the ket in (15) is replaced with the state
$\lvert\psi_{\times}\rangle=\lvert\underbrace{\downarrow\downarrow\dots\downarrow}_{N-1}\uparrow,\downarrow,\uparrow\dots\uparrow\rangle.$ (27)
Following Pérez-García et al. (2022), we obtain that the resulting coupling-averaged probability, divided by the coupling-averaged Loschmidt echo, equals $\mathcal{P}$. We elaborate further on this aspect in §B.1.
In summary, the expectation value of a Polyakov loop can be analysed on the spin chain by testing how the results vary when a simple impurity is introduced into the initial state.
Even without any reference to the gauge theory, we predict the ratio of the Loschmidt echoes with and without impurities to be an order parameter discerning between a phase with low disorder, $0<a\ll 1$, and a phase of large disorder, thus large entropy, $a\gg 1$. This is a novel claim for the coupling averaged Heisenberg XX chain and may be of independent interest. To quantitatively test this prediction with the current state of the art technologies seems an interesting avenue to pursue.
Improved estimates of the entropy
@@
3. The spin chain can reproduce the perturbative corrections in the ’t Hooft coupling to the BH entropy (§B.4). Surprisingly, this major conceptual step requires no further sophistication of the experimental setup. These corrections are incorporated without complicating the Hamiltonian nor the initial state. The implementation simply requires that the random coupling is sampled with random standard deviation — see §B.4 for details. The BH entropy is computed by ‘nested’ averages of the Loschmidt echo.
@@ IV Discussion
Assuming the AdS5/CFT4 correspondence, we showed that the Hawking–Page transition with maximal supersymmetry is reproducible on an elementary 1d spin chain. The result is based on the mathematical identity (26), relating the entropy to the Loschmidt echo of the XX chain with an average over the randomized coupling. We also studied order operators for this transition. Additionally, we enriched the chain in various ways to simulate more realistic features of the BH entropy.
Unlike the planar integrability of $\mathcal{N}=4$ SYM (reviewed in §C), the XX chain used here is an auxiliary device manifesting the phase transition, rather than being a subsector of SYM. We map one specific observable to the averaged thermal Loschmidt echo in the spin chain.
The outcome highlights a universality property: introducing entropy in the 1d chain through the average, it emulates the behaviour of (superficially unrelated) BH systems in 4+1 dimensions.
It is natural to inquire about the extent to which the spin chain captures structural properties of BH physics, besides the jump in entropy at the Hawking–Page transition. In §D, we elaborate on two concrete avenues for future research, aimed at characterizing the von Neumann algebra and the microstates of the BH via spin chain techniques.
@@ Acknowledgements.
We are grateful Hamed Adami, Elliott Gesteau and especially Sofyan Iblisdir for discussions. LS thanks the Departamento de Análisis Matemático y Matemática Aplicada, Universidad Complutense de Madrid for warm hospitality at various stages of this project. The work has been financially supported by Comunidad de Madrid (grant QUITEMAD-CM, ref. P2018/TCS-4342), by Universidad Complutense de Madrid (grant FEI-EU-22-06), by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2025 Agenda. DPG also acknowledges financial support from the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (“Severo Ochoa Programme for Centres of Excellence in R&D” CEX2019- 000904-S and grant PID2020-113523GB-I00) as well as from CSIC Quantum Technologies Platform PTI-001. The work of LS is supported by the Shuimu Scholars program of Tsinghua University and by the Beijing Natural Science Foundation project IS23008 (“Exact results in algebraic geometry from supersymmetric field theory”).
@@
Our study of the AdS5 entropy across the Hawking–Page transition is done in two steps: first assume the validity of the AdS/CFT correspondence, which maps the supergravity problem to a four-dimensional gauge theory problem, and then take the large $N$ limit in the gauge theory. We now present the pertinent quantities directly in $\mathcal{N}=4$ SYM.
@@ A.2 Thermal partition function
The entropy in AdS5 is predicted to be dual to the thermal partition function, and we have computed
@@
$S=\log(\text{\# microstates}).$ (32)
An alternative take on the relation (31) uses that, in the BH phase $T>T_{\mathrm{HP}}$, the near-horizon geometry is well-known to be that of an AdS2, which implies the dual theory to reduce to an effective quantum mechanics along a thermal circle. This effective field theory approach lies at the heart of the derivation of (6), as we now review.
@@ A.4 Weak coupling corrections and deconfinement
In summary, the simplified partition function (8) captures all the distinctive properties of the Hawking–Page transition, and a more realistic approximation, which includes a coexistence region for thermal AdS and BH, is described by the improved formula (33).
@@
B.1 Polyakov loop
From the point of view of the gauge theory, the Hawking–Page transition is holographically dual to a deconfinement transition at $T=T_{\rm HP}$ Witten (1998b). Famously, a probe of deconfinement ’t Hooft (1978) is provided by the so-called Polyakov loop Polyakov (1978), placed along the circle $\mathbb{S}^{1}$ of radius $1/T$ in this context. Let us denote $\mathcal{P}$ the Polyakov loop expectation value, and let $A$ be the $SU(N)$ or $U(N)$ gauge field. Then,
where $\langle\cdot\rangle_{\mathcal{N}=4}$ stands for the correlation function in the full-fledged $\mathcal{N}=4$ CFT. Wilson loops in general, and $\mathcal{P}$ in particular, are charged under the centre symmetry of $\mathcal{N}=4$ SYM theory. This is directly seen by the fact that (34) is rotated by a phase when the gauge field $A$ is shifted by a flat 1-form in the centre of the gauge group. This makes $\mathcal{P}$ an order parameter for the deconfinement transition in $\mathcal{N}=4$ SYM: it must vanish in the low temperature phase, in which the centre symmetry is preserved, and acquires a vacuum expectation value in the high temperature phase, in which deconfinement spontaneously breaks the centre symmetry.
$\mathcal{P}=\left\langle\mathrm{Tr}U\right\rangle$ (35)
[…]
$\exp\left(N\varepsilon\mathrm{Tr}U\right)$ (36)
and, denoting $\hat{S}(\varepsilon)$ the entropy modified in this way,
On the spin chain side, it was shown in Pérez-García et al. (2022) that matrix model averages such as (35) correspond to impurities in the preparation of the initial state $\lvert\psi_{0}\rangle$. Then, $\mathcal{P}$ corresponds to the ratio of amplitudes with and without impurity. More precisely, let $\lvert\psi_{\times}\rangle$ be as in (27), and denote $\mathcal{G}_{N}^{\times}(J)$ (respectively $\hat{\mathcal{G}}_{N}^{\times}(J)$) the analogue of $\mathcal{G}_{N}(J)$ (resp. $\hat{\mathcal{G}}_{N}(J)$) constructed using $\lvert\psi_{\times}\rangle$ as initial state, instead of $\lvert\psi_{0}\rangle$. Explicitly,
exactly as in (26). The centre symmetry of $\mathcal{N}=4$ SYM does act on the spin chain with the defect, rotating the amplitude $\hat{\mathcal{G}}_{N}^{\times}$ by a phase. However, any physical implementation would automatically select a real saddle point for $\mathcal{P}$, since the hypothetical experiment would only be measuring the real probability (left invariant by the action of the centre symmetry) and not the amplitude. Stated differently, any concrete implementation of (40) would have built in the analogue of the regulating procedure described in Aharony et al. (2004) to obtain $\mathcal{P}\neq 0$.
It is also noteworthy that the definition of $\mathcal{P}$ can be modified in both $\mathcal{N}=4$ SYM and the spin chain. On the field theory side, it is possible to examine loops in higher-dimensional representations of the gauge group, such as the rank-$p$ antisymmetric representation. On the spin chain side, a feasible approach is to design an initial state $\lvert\psi_{\times}\rangle$ where the final spin $\downarrow$ is shifted $p\geq 1$ locations away from the other $N-1$ spins $\downarrow$. Notwithstanding this alteration, Identity (40) remains valid.
B.2 Improving the estimate of the BH entropy